A tungsten wire has a radius of 0.076 mm and is heated from 20.0 to 1305 °C. The temperature coefficient of resistivity is α = 4.5 10-3 (C°)−1. When 120 V is applied across the ends of the hot wire, a current of 1.9 A is produced. How long is the wire? Neglect any effects due to thermal expansion of the wire.
Answer #1

x r = 0.076mm AT = (1305- 20) = 1285 °C 4.5 x 10-30-1 V = 220v ILGA R = V - 120 - 63. 1bir I 7-9 R= p I - 63.16 Af=&POT P = S